OpenStudy (im_an_angel):

Need some help.... The length of a rectangle is 6 inches more than its width, and its perimeter is 32 inches. Find the width of the rectangle. A. 5 B. 11 C. 16

OpenStudy (anonymous):

Perimeter is just the sum of all the sides. If a rectangle has four sides, & two pairs of sides that are the same length, then it will be easier to find this. Basically, 2(L+W)=2(6+x)= 12+2x.

OpenStudy (anonymous):

Subtract 32-12. Then divide your answer by two.

OpenStudy (im_an_angel):

Ok I get it now. Thanks a bunch :D @ Brostep0s

OpenStudy (im_an_angel):

@Brostep0s